OpenStudy (anonymous):

Explain two ways you could evaluate (f + g)(2)

OpenStudy (ash2326):

Supposedly we have \[f(x)=x\] and \[g(x)=x^2\] \[(f+g)(x)=x+x^2\] so if we need to find \[(f+g)(2)\] we could first find \((f+g)(x)\) and put x=2 or we could find f(2) and g(2) and add them

OpenStudy (ash2326):

@llalaa do you get this?

OpenStudy (anonymous):

Yes, thank you(: